OpenStudy (emilyjones284):

MEDAL AND FAN! (: A set of data with a mean of 45 and a standard deviation of 8.3 is normally distributed. Find the value that is –2 standard deviations away from the mean.

jimthompson5910 (jim_thompson5910):

z = (x - mu)/sigma -2 = (x - 45)/8.3 solve for x

OpenStudy (emilyjones284):

What would be the first step to solving for x?

jimthompson5910 (jim_thompson5910):

multiplying both sides by 8.3

jimthompson5910 (jim_thompson5910):

-2 = (x - 45)/8.3 -2*8.3 = x - 45 -16.6 = x - 45

OpenStudy (emilyjones284):

so then its -16.6(x-45)?

jimthompson5910 (jim_thompson5910):

close, look at what I wrote above

OpenStudy (emilyjones284):

Ohhhhhh okay so its 28.4?

OpenStudy (emilyjones284):

wait thats not an option for an answer

jimthompson5910 (jim_thompson5910):

what are your options?

OpenStudy (emilyjones284):

20.1, 78.9, 69.9, 48.0

jimthompson5910 (jim_thompson5910):

hmm I'm getting 28.4 as well, so not sure why it's not showing up

OpenStudy (emilyjones284):

did we do it wrong?

jimthompson5910 (jim_thompson5910):

double check all the numbers to make sure they are correct (like the 45 or 8.3 or -2)

jimthompson5910 (jim_thompson5910):

no we did it correctly

OpenStudy (emilyjones284):

Oh wow I'm sorry instead of -2 it is +3, I dont know how I did that

jimthompson5910 (jim_thompson5910):

that's ok, glad you checked

jimthompson5910 (jim_thompson5910):

so you need to solve 3 = (x - 45)/8.3

OpenStudy (emilyjones284):

Okay I got 69.9 (:

jimthompson5910 (jim_thompson5910):

me too

OpenStudy (emilyjones284):

Thanks (:

jimthompson5910 (jim_thompson5910):

yw